# Arrow-Debreu Meets Kyle: Price Discovery Across Derivatives

### v39

@@ 1 Introduction
While cross-price impact is the local (marginal) footprint of learning from order flow, the corresponding global object is the joint system of supply schedules for all AD securities—an asymmetric-information pricing kernel that prices all contingent-claims, and hence derivatives. [^1] This pricing kernel is a high-dimensional, nonlinear object. It maps cross-market order-flow to the cross-section of state prices; its marginal quantity—price impact—is already an infinite-dimensional linear operator. The extensive derivative-pricing literature since Black and Scholes (1973) has largely focused on hedging and replication under complete information. Yet hedging is clearly not the only motive for trading derivatives. We show how informed trading explains observed price behavior that hedging does not account for. In particular, the asymmetric-information pricing kernel generates the volatility smile across option strikes, a well-documented empirical regularity of option-implied volatilities. Reduced-form frameworks descended from Black-Scholes can be tuned to match the smile, but they effectively treat this market outcome—an object implied by option prices—as an input to be fitted. [^2] A large financial econometrics literature fits the volatility smile/surface by modifying components of Black-Scholes-type models. Prominent approaches include the jump-diffusion models (Merton (1976), Bates (1991)), stochastic volatility models (Heston (1993), Bates (1996), Duffie et al. (2000), Britten-Jones and Neuberger (2000), Bates (2000), Aït-Sahalia and Kimmel (2007), Christoffersen et al. (2010), Aït-Sahalia et al. (2021)), and local volatility models (Dupire (1994), Berestycki et al. (2002), and Carr and Cousot (2012)).
Our model also yields a simple, practical expression of equilibrium informed demand. The informed trader trades on his private signal via a long-short contingent-claim portfolio: he goes long the signal-implied payoff distribution and shorts those implied by alternative signals. In the options specialization, this translates to concrete multi-leg trades that target the chosen aspects of the underlying payoff distribution; for example, volatility information induces the straddle, a workhorse volatility strategy. More broadly, equilibrium informed demand rationalizes familiar, long-standing option strategies—going beyond the existing (mostly single-contract) models that are silent on observed multi-option trading practices. This equilibrium portfolio characterization provides a constructive guide for designing, and empirically interpreting, option portfolios that load on particular informational dimensions
@@
Kyle (1985) is the seminal framework for single-asset price discovery. Subsequent work extends this framework to settings with multiple insiders and information dispersed across securities—e.g., Holden and Subrahmanyam (1992); Caballe and Krishnan (1994); Foster and Viswanathan (1996); Back et al. (2000). All of these models impose a linear-Gaussian structure—joint normality of payoffs, signals, and noise—leaving price discovery across claims with nonlinear payoffs beyond their scope. Rochet and Vila (1994) relaxes normality but remains in a single-asset setting. We carry the Kyle price-discovery logic to Arrow-Debreu markets. The Arrow-Debreu pricing kernel is generically infinite-dimensional—even with finitely many primitive assets—whereas the linear-Gaussian setting is finite-dimensional by construction. Our departure from that setting is therefore both conceptual and methodological, not merely a larger asset menu.
@@
Despite this robust evidence of joint price formation across option contracts, existing informed-trading models (e.g., Back (1993); Biais and Hillion (1994); Brennan and Cao (1996); Easley et al. (1998); Collin-Dufresne et al. (2021)) remain confined to single-option settings and therefore cannot speak to the cross-contract mechanisms at the core of these empirical findings—cross-strike dynamics, implied-volatility-surface formation, and multi-option trades such as the straddle. Our model yields cross-option equilibrium restrictions that sharpen and generalize the empirical intuition while generating additional testable implications. For instance, we identify cross-strike volume imbalance as the equilibrium channel behind the return predictability of signed option volume documented by Pan and Poteshman (2006), and provide the first equilibrium rationale for the volatility-forecasting power of straddle demand observed in Ni et al. (2008) and Goyal and Saretto (2009).
There have also been theoretical studies of informed trading across multiple securities in competitive (non-strategic) settings, such as Admati (1985), Malamud (2015), and Chabakauri et al. (2022). Without strategic order flow, these models cannot generate the derivatives trading strategies used in practice. [^3] This mirrors the familiar distinction between Kyle (1985) and Grossman and Stiglitz (1980) settings in the single-asset case. Similarly, the insight mapping payoff characteristics to observable cross-price impact and cross-strike regularities is possible only in a strategic setting. Moreover, Admati (1985) and Chabakauri et al. (2022) impose restrictive parametric assumptions on the payoff distribution, while Malamud (2015) relies on properties intrinsic to the continuum. By contrast, we impose no restrictions on the payoff distribution, and our results do not depend on whether the state space is discrete or continuous.
@@
The reduction of Section 6 is the key modeling innovation that delivers tractability at our full Arrow-Debreu generality and yields an equilibrium characterized by a single endogenous constant. Readers who prefer the high-level overview without the formal developments can see the reduction distilled in Section 6.3, which links directly back to the simplified environment of Section 2.
@@ 2 Basic Intuition
Cross price impact is positive when two securities are informational complements: learning that one security has a high payoff raises the insider’s belief that the other security also pays. Securities $x_{1}$ and $x_{3}$ are complements: their payoff covariance is $\tfrac{1}{8}$, so their cross price impact is positive. By contrast, the other two pairs have payoff covariance $\tfrac{1}{8}$, resulting in negative cross price impact. Payoff covariances determine the extent to which order flow for one security reveals information about others, and thus determine price impact across markets. This intuition extends directly to the general model developed in subsequent sections.
@@
In choosing the scale $\alpha$ of his long-short portfolio, the insider trades off the marginal payoff from exploiting the true payoff distribution against the marginal cost of trading. That cost has two components. First, there is the cost of executing at prevailing prices, absent any price impact. Second, there is an informational component: larger trades make order flow more informative for the market maker, which amplifies cross-market price impact and erodes the insider’s informational advantage. The optimal scale $\alpha^{*}$ balances these margins (see Figure 1a). Equilibrium is characterized by this single endogenous constant—nontrivially, this same equilibrium structure carries over to the general model (see Section 6.3). Allowing correlated noise trades, varying noise intensity, and non-uniform priors has no material effect on the main results. [^7] Correlated Noise Trades Allowing noise trades to be correlated (holding their marginal variances fixed) changes the pattern of uninformed order flow, but not equilibrium price impact at the margin. The reason is that the market maker conditions on the noise covariance when extracting information, so predictable co-movement in noise is effectively filtered out. For example, if noise co-moves in the opposite direction of payoff co-movement, the insider optimally trades more aggressively because order flow is better camouflaged; anticipating this, the market maker adjusts his supply schedules, leaving the marginal price response (the impact matrix) unchanged. This is analogous to the single-asset case where biased noise shifts levels but not the slope of price impact.  [^8] Varying Noise Intensity Suppose noise intensity differs across markets, with $\epsilon_{i}\!\stackrel{{\scriptstyle d}}{{\sim}}\!\mathcal{N}(0,\sigma_{i}^{2})$ for $i=1,2,3$. Then the market maker “precision-weights” order flow, and the sufficient statistic becomes $\Delta(\omega)=\frac{\omega_{1}}{\sigma_{1}^{2}}+\frac{\omega_{3}}{\sigma_{3}^{2}}-\frac{\omega_{2}}{\sigma_{2}^{2}}.$ This is intuitive. A larger $\sigma_{i}$ means order flow in market $i$ is noisier (lower signal-to-noise), so a given trade there is less informative and receives less weight in inference. Consequently, higher noise trading intensity attenuates price impact both within a market and across markets.  [^9] General Prior With a general prior $(\pi_{0}(s_{1}),\pi_{0}(s_{2}))$, the posterior odds are just the likelihood ratio from (1) reweighted by the prior odds $\frac{\pi_{1}(s_{1}\mid\omega)}{\pi_{1}(s_{2}\mid\omega)}=\exp\!\big(2\alpha\,\Delta(\omega)\big)\cdot\frac{\pi_{0}(s_{1})}{\pi_{0}(s_{2})},\qquad\Delta(\omega)=\omega_{1}+\omega_{3}-\omega_{2}.$ Intuitively, a higher $\pi_{0}(s_{1})$ means the market maker needs less order-imbalance evidence to attribute trading to the $s_{1}$ “story.” In the toy volatility interpretation, co-buying in the two straddle legs is therefore taken as stronger evidence of high volatility, which amplifies their put-call cross price impact.
@@
Under the high volatility signal, the insider therefore loads on both put and call, so buy pressure in either leg is interpreted as evidence favoring high volatility, which pushes up the other leg as well. Consequently, the cross price impact between the put-call pair is informative about future volatility. In the general model, we systematically extend this insight to actual options: cross-option price impact reveals information about the underlying payoff distribution, yielding empirical predictions for how option prices and implied-volatilities adjust as price discovery unfolds.
@@ 3 General Model
Assume the asset value follows a normal prior distribution $\pi_{0}\!\stackrel{{\scriptstyle d}}{{\sim}}\!\mathcal{N}(v_{0},\sigma_{v}^{2})$ with prior mean $v_{0}$ and variance $\sigma_{v}^{2}$, and that noise trades follow $\varepsilon\!\stackrel{{\scriptstyle d}}{{\sim}}\!\mathcal{N}(0,\sigma_{\varepsilon}^{2})$. Under these assumptions, there exists a linear equilibrium
@@
Remark 3.2.
The general economic intuition we bring out is robust to whether the state and signal spaces are discrete or continuous. A continuous setup—such as the normal prior/payoff environment of Kyle (1985)—can be approximated arbitrarily well by a discrete signal-state grid. We consider the discrete model here to focus on the economic results and minimize technicality. A fully rigorous continuous-space treatment (allowing general priors and signal structures) is provided in Keller and Tseng (2026).
@@ 4.1 The Posterior and Pricing Kernel
that is, how closely the realized order flow aligns with the expected signal-contingent demand profiles. These scores therefore form the sufficient statistic for his inference.
@@ 4.2 Sufficient Statistic
Theorem 4.1 makes this intuition precise under arbitrary information structures. For each signal $s_{k}$, the market maker computes a precision-weighted alignment score
@@
His posterior is therefore determined by the vector of alignment scores (11)—one for each information type. The weights $1/\sigma_{i}^{2}$ are the signal-to-noise adjustment: trades in noisier markets (larger $\sigma_{i}$) are less informative and therefore receive less weight in inference. At the extreme, if $\sigma_{i}\to\infty$ for all $i$, order flow becomes pure noise, the alignment scores shrink to zero, and the posterior reduces to the prior.
@@
We will refer to (12) as the market maker’s alignment score for signal $s_{k}$. We will define the analogous impact exposure when $\omega$ is replaced by a candidate insider portfolio $W$, since that quantity will govern the portfolio’s price impact in Section 5.
@@
(i) The precision-weighted alignment score of order flow $\omega$ with $\widetilde{W}(\,\cdot\,|s_{k})$ is
@@
(ii) The precision-weighted impact exposure of an insider portfolio $W(\,\cdot\,)$ to $\widetilde{W}(\,\cdot\,|s_{k})$ is
@@ 6 Equilibrium
For expositional convenience, we maintain Assumption 1 in this section. It is without loss of generality and serves only to streamline the derivations. Starting from any prior $\pi_{0}$, we can rescale payoffs and work as if the prior were uniform on $S$ by replacing $\eta(\,\cdot\,|s_{k})$ with $\frac{\pi_{0}(s_{k})}{1/|S|}\,\eta(\,\cdot\,|s_{k})$ for each $s_{k}\in S$. Once we characterize equilibrium under the uniform prior, comparative statics for arbitrary priors follow immediately. For instance, if two securities both yield higher payoffs under some $s_{k}$ compared to the other signals, then their cross price impact increases with the prior probability of $s_{k}$. For instance, if two securities tend to pay off together under some $s_{k}$, then putting more prior weight on $s_{k}$ makes their co-movement in order flow more diagnostic, so their equilibrium cross price impact increases with $\pi_{0}(s_{k})$. This generalizes from the toy model of Section 2; see footnote 9.
@@ 6.1 Canonical Game
The information intensity matrix $\bf L$} is the cross-market analogue of the signal-to-noise ratio $\sigma_{v}/\sigma_{\varepsilon}$ from the single-asset Kyle setting (6). It aggregates the (noise-adjusted) co-variations of the signal-implied payoff profiles $\{\eta(\,\cdot\,\mid s_{k})\}_{k}$ across claims, so a larger $\bf L$} means that cross-market order flow is more diagnostic of which information story is active. [^16] Here, “larger” is in the positive-semidefinite order. This order is partial—unlike the single-asset setting, information intensity is not always directly comparable between two specifications.
@@
so the $k$th column $\beta^{(k)}$ collects the insider’s equilibrium loadings across payoff distributions when signal $s_{k}$ is realized. The corresponding informed demand is then
@@ 6.2 Symmetry and One-Parameter Equilibrium
Imposing insider optimality within this symmetric class reduces the fixed point condition to a single scalar equation in $\alpha$. Theorem 6.3 shows that it admits a solution $\alpha^{*}>0$, delivering equilibrium strategy matrix $D^{*}=\alpha^{*}{\bf Q}$.
@@
The intuition underlying Theorem 6.3 is the same as that highlighted in the toy setting of Section 2, now shown to be robust in full Arrow-Debreu generality. Raising $\alpha$ scales up the long-short portfolio and increases the insider’s exposure to the true story, but it also makes aggregate order flow more informative, which erodes the insider’s informational edge. Equilibrium $\alpha^{*}$ obtains at the knife-edge where this tradeoff balances.

### v40

@@ 1 Introduction
While cross-price impact is the local (marginal) footprint of learning from order flow, the corresponding global object is the joint system of supply schedules for all AD securities—an asymmetric-information pricing kernel that prices all contingent-claims, and hence derivatives. [^1] This pricing kernel is a high-dimensional, nonlinear object. It maps cross-market order-flow to the cross-section of state prices; its marginal quantity—price impact—is already an infinite-dimensional linear operator. The extensive derivative-pricing literature since Black and Scholes (1973) has largely focused on hedging and replication under complete information. Yet hedging is clearly not the only motive for trading derivatives. We show how informed trading explains observed price behavior that hedging alone does not account for. In particular, the asymmetric-information pricing kernel endogenously generates the volatility smile across option strikes, a well-documented empirical regularity of option-implied volatilities. By contrast, reduced-form frameworks descended from Black–Scholes can only fit the smile by ex post calibration, an exercise that offers no structural explanation—it treats a market outcome implied by option prices as an exogenous input to be fitted. [^2] A large financial econometrics literature fits the volatility smile/surface by modifying components of Black-Scholes-type models. Prominent approaches include the jump-diffusion models (Merton (1976), Bates (1991)), stochastic volatility models (Heston (1993), Bates (1996), Duffie et al. (2000), Britten-Jones and Neuberger (2000), Bates (2000), Aït-Sahalia and Kimmel (2007), Christoffersen et al. (2010), Aït-Sahalia et al. (2021)), and local volatility models (Dupire (1994), Berestycki et al. (2002), and Carr and Cousot (2012)).
Our model also yields a simple, practical expression of equilibrium informed demand. The informed trader trades on his private signal via a long-short contingent-claim portfolio: he goes long the signal-implied payoff distribution and shorts those implied by alternative signals. In the options specialization, this translates to concrete multi-leg trades that target the chosen aspects of the underlying payoff distribution; for example, volatility information induces the straddle, a workhorse volatility strategy. More broadly, equilibrium informed demand rationalizes familiar, long-standing option strategies—going beyond the existing (mostly single-contract) models that are silent on observed multi-option trading practices. This equilibrium portfolio characterization provides a constructive guide for designing, and empirically interpreting, option portfolios that load on particular informational dimensions.
@@
Kyle (1985) is the seminal framework for single-asset price discovery. Subsequent work extends this framework to settings with multiple insiders and information dispersed across securities—e.g., Holden and Subrahmanyam (1992); Caballe and Krishnan (1994); Foster and Viswanathan (1996); Back et al. (2000). All of these models impose a linear-Gaussian structure—joint normality of payoffs and signals—leaving price discovery across claims with nonlinear payoffs beyond their scope. Rochet and Vila (1994) relaxes normality but remains in a single-asset setting. We carry the Kyle price-discovery logic to Arrow-Debreu markets. The Arrow-Debreu pricing kernel is generically infinite-dimensional—even with finitely many primitive assets—whereas the linear-Gaussian setting is finite-dimensional by construction. Accordingly, our departure from that setting calls for both new conceptual insights and methodological tools.
@@
Despite this robust evidence of joint price formation across option contracts, existing informed-trading models (e.g., Back (1993); Biais and Hillion (1994); Brennan and Cao (1996); Easley et al. (1998); Collin-Dufresne et al. (2021)) remain confined to single-option settings and therefore cannot speak to the cross-contract mechanisms at the core of these empirical findings—cross-strike dynamics, implied-volatility-surface formation, and multi-option trades such as the straddle. Our model yields cross-option equilibrium restrictions that consolidate and sharpen this body of evidence while generating additional testable predictions. For instance, we identify cross-strike volume imbalance as the equilibrium channel behind the return predictability of signed option volume documented by Pan and Poteshman (2006), and provide the first equilibrium rationale for the volatility-forecasting power of straddle demand observed in Ni et al. (2008) and Goyal and Saretto (2009).
There have also been theoretical studies of informed trading across multiple securities in competitive (non-strategic) settings, such as Admati (1985), Malamud (2015), and Chabakauri et al. (2022). Without strategic order flow, these models cannot generate the derivatives trading strategies used in practice. [^3] This mirrors the familiar distinction between Kyle (1985) and Grossman and Stiglitz (1980) settings in the single-asset case. Similarly, the insight mapping payoff characteristics to observable cross-price impact and cross-strike regularities is possible only in a strategic setting. Moreover, Admati (1985) and Chabakauri et al. (2022) impose restrictive parametric assumptions on the payoff distribution, while Malamud (2015) relies on properties intrinsic to the continuum. By contrast, we impose no restrictions on the payoff distribution, and our results do not depend on the discrete–continuous modeling choice.
To foreground the economic intuition, we adopt the discrete formulation in this paper. At the level of intuition, a continuous setup—such as the Gaussian prior-payoff environment of Kyle (1985)—is well approximated by a discrete signal-state grid. A rigorous treatment of the continuous formulation is provided in the companion paper Keller and Tseng (2026), which addresses the substantive technical considerations it entails.
@@
The reduction of Section 6 is the key insight that renders transparent the extremely tight one-dimensional restriction governing equilibrium under full Arrow–Debreu generality. Readers who prefer the high-level overview without the formal developments can find the reduction distilled in Section 6.3, which links directly back to the simplified environment of Section 2.
@@ 2 Basic Intuition
Cross price impact is positive when two securities are informational complements: learning that one security has a high payoff raises the insider’s belief that the other security also pays. Securities $x_{1}$ and $x_{3}$ are complements: their payoff covariance is $\tfrac{1}{8}$, so their cross price impact is positive. By contrast, the other two pairs have payoff covariance $\tfrac{1}{8}$, resulting in negative cross price impact. Payoff covariances determine the extent to which order flow for one security reveals information about others, and thus determine price impact across markets. In the sections that follow, we bring out this intuition under the general model.
@@
In choosing the scale $\alpha$ of his long-short portfolio, the insider trades off the marginal gains from exploiting the true payoff distribution against the marginal trading cost. That cost has two components. First, there is the cost of transacting at prevailing prices, absent any price impact. Second, there is an informational component: larger trades make order flow more informative for the market maker, which steepens cross-market price impact and erodes the insider’s information rent. The optimal trading scale $\alpha^{*}$ balances these margins (see Figure 1b). Equilibrium is pinned down by this single endogenous constant $\alpha^{*}$—and, nontrivially, the same one-parameter structure carries over to the general model (see Section 6.3 for intuition). Allowing correlated noise trades, varying noise intensity, and non-uniform priors has no material effect on the core mechanisms, either in this simplified setting or in the general model. [^7] Correlated Noise Trades Suppose noise trades are correlated (holding their marginal variances fixed). This changes the pattern of uninformed flow across claims but not marginal price impact. This is because the market maker filters out the predictable noise correlation and prices only the residual surprise. For example, if noise tends to co-move opposite to payoff co-movement, the insider would hide in the offsetting flow and trade more aggressively; anticipating this, the market maker shifts supply schedules but the marginal price responses are unchanged. This mirrors the single-asset case, where biased noise shifts the level of the supply schedule but not Kyle’s lambda.  [^8] Varying Noise Intensity Suppose noise intensity differs across markets, with $\epsilon_{i}\sim\mathcal{N}(0,\sigma_{i}^{2})$ for $i=1,2,3$. Then the market maker’s order-imbalance sufficient statistic becomes $\Delta(\omega)=\frac{\omega_{1}}{\sigma_{1}^{2}}+\frac{\omega_{3}}{\sigma_{3}^{2}}-\frac{\omega_{2}}{\sigma_{2}^{2}}.$ This is the standard signal-to-noise intuition—flow in a noisier market (larger $\sigma_{i}$) receives less weight in inference. Therefore, trades in noisier contracts have lower price impact both within, and across markets. 99footnotemark: 9
@@
Under the high volatility signal, the insider therefore loads on both put and call, so buy pressure in either leg is interpreted as evidence favoring high volatility, which pushes up the other leg as well. Consequently, the cross price impact between the put-call pair is informative about future volatility. In the general model, we systematically extend this insight to actual options: cross-option price impact reveals information about the underlying payoff distribution, yielding empirical predictions for how option prices and implied-volatilities adjust as price discovery unfolds (see Section 8).
@@ 3 General Model
Assume the asset value follows a normal prior distribution $\pi_{0}\!\stackrel{{\scriptstyle d}}{{\sim}}\!\mathcal{N}(v_{0},\sigma_{v}^{2})$ with prior mean $v_{0}$ and variance $\sigma_{v}^{2}$, and that noise trades follow $\varepsilon\!\stackrel{{\scriptstyle d}}{{\sim}}\!\mathcal{N}(0,\sigma_{\varepsilon}^{2})$. (As noted in the introduction, our results are fully robust to whether the signal and state spaces are discrete or continuous—intuitively, this joint normal setup can be approximated by a discrete grid.) Under these assumptions, there exists a linear equilibrium
@@ 4.1 The Posterior and Pricing Kernel
which measure how closely the realized order flow aligns with the expected signal-contingent demand profiles. These scores therefore form the sufficient statistic for his inference.
@@ 4.2 Sufficient Statistic
Theorem 4.1 makes this intuition precise under arbitrary information structures. For each signal $s_{k}$, the market maker computes a noise-adjusted alignment score
@@
His posterior is therefore determined by the profile of alignment scores (11)—one for each information type. The weights $1/\sigma_{i}^{2}$ are the signal-to-noise adjustment: trades in noisier markets (larger $\sigma_{i}$) are less informative and therefore receive less weight in inference. At the extreme, if $\sigma_{i}\to\infty$ for all $i$, order flow becomes pure noise, the alignment scores shrink to zero, and the posterior reduces to the prior.
@@
We also define the analogous impact exposure by replacing order flow $\omega$ with a candidate insider portfolio $W$ in the alignment score (12).
@@
(i) The noise-adjusted alignment score of order flow $\omega$ with $\widetilde{W}(\,\cdot\,|s_{k})$ is
@@
(ii) The noise-adjusted impact exposure of an insider portfolio $W(\,\cdot\,)$ to $\widetilde{W}(\,\cdot\,|s_{k})$ is
@@ 6 Equilibrium
For expositional convenience, we maintain Assumption 1 in this section. It is without loss of generality and serves only to streamline the derivations. Starting from any prior $\pi_{0}$, we can rescale payoffs and work as if the prior were uniform on $S$ by replacing $\eta(\,\cdot\,|s_{k})$ with $\frac{\pi_{0}(s_{k})}{1/|S|}\,\eta(\,\cdot\,|s_{k})$ for each $s_{k}\in S$. Once we characterize equilibrium under the uniform prior, comparative statics for arbitrary priors follow immediately. For instance, if two securities both yield higher payoffs under some $s_{k}$ compared to the other signals, then their cross price impact increases with the prior probability of $s_{k}$. For instance, if two securities tend to pay off together under some $s_{k}$, then putting more prior weight on $s_{k}$ makes their co-movement in order flow more diagnostic, so their equilibrium cross price impact increases with $\pi_{0}(s_{k})$. This generalizes from the toy model of Section 2; see footnote LABEL:fn:_general_prior.
@@ 6.1 Canonical Game
The information intensity matrix $\bf L$} is the cross-market analogue of the information intensity parameter $\sigma_{v}/\sigma_{\varepsilon}$ from the single-asset Kyle setting (6). It aggregates the (noise-adjusted) co-variations of the signal-implied payoff profiles $\{\eta(\,\cdot\,\mid s_{k})\}_{k}$ across claims, so a larger $\bf L$} means that cross-market order flow is more diagnostic of which information story is active. [^16] Here, “larger” is in the positive-semidefinite order. This order is partial—unlike the single-asset setting, information intensity is not always directly comparable between two specifications.
@@
so the $k$-th column $\beta^{(k)}$ collects the insider’s equilibrium loadings across payoff distributions when signal $s_{k}$ is realized. The corresponding informed demand is then
@@ 6.2 One-Parameter Equilibrium
Imposing insider optimality within this symmetric class reduces the fixed point condition to a single scalar equation in $\alpha$. Theorem 6.3 shows that it admits a solution $\alpha^{*}>0$, delivering equilibrium strategy matrix $D^{*}=\alpha^{*}{\bf Q}$. Thus, $\alpha^{*}$ is the sole endogenous constant characterizing equilibrium.
@@
The intuition underlying Theorem 6.3 is the same as that highlighted in the toy setting of Section 2, now shown to be robust in full Arrow-Debreu generality. Raising $\alpha$ scales up the long-short portfolio and increases the insider’s exposure to the true story, but it also makes aggregate order flow more informative, which erodes the insider’s marginal information rent. Equilibrium $\alpha^{*}$ obtains at the knife-edge where this tradeoff balances.